OpenStudy (chycora):

Most flat screen TVs today are made with an aspect ratio (ratio of the length to the height) of 16:9. Xenock Electronics' top selling TV is their 28-inch model, whose viewing area has a length of 24 inches and a height of 13.5 inches. This TV costs $226.99. They have three main competitors that claim to offer a better deal per viewing area.

OpenStudy (chycora):

Phil's Supercenter claims their 18-inch model, whose viewing area length is \[\frac{ 2 }{ 3 }\] of the length of Xenock's 28-inch model, is a better deal per viewing area at $152.99. Tyler's Electronics claims their 33-inch model, whose viewing area height is 20% taller than Xenock's 28-inch model, is a better deal per viewing area at $276.99. Beltre's Superstore claims their 82-inch model, whose viewing area dimensions are 300% as large as Xenock's 28-inch model, is a better deal per viewing area at $2,079.99. List the TV's in order of best deal per viewing area. Assume all TV's have the same aspect ratio.

OpenStudy (anonymous):

nope no smarts gang

OpenStudy (chycora):

wat

OpenStudy (anonymous):

nun

OpenStudy (anonymous):

who the cat

OpenStudy (chycora):

idk pro somebody who was goin 2 help me cuz i need help

OpenStudy (chycora):

@iGreen

OpenStudy (chycora):

@tkhunny

OpenStudy (chycora):

can u guys plz help me

OpenStudy (tkhunny):

The hard part is assuming they all have the same aspect ratio. How tall is Xenock's 28? Let's call it T. How wide is Xenock's 28? Let's call it W. Then \(T^{2} + W^{2} = 28^{2}\) If Phil's 18 has 2/3 the width as Xenock, it must also have 2/3 the height or the aspect ration changes. Then Phil's \([(2/3)T]^{2} + [(2/3)W]^{2} = 18^2\) Or \(T^{2} + W^{2} = [18(3/2)]^2 = [9*3]^2 = 27^2\) And suddenly, we are not making any sense. Well, I suggest you play with the other sets and see if something useful falls out. Phil and Xenock don't seem to be making much sense.